# Bridging the Gap: OPTICAM Reveals the Hidden Spin of the WZ Sge Star GOTO 065054.49+593624.51

### v1

@@ 1 Introduction
Cataclysmic variable stars (CVs) are binary systems in which a low-mass secondary star, filling its Roche lobe, transfers mass to a white dwarf (WD) primary. In systems where the magnetic field is dynamically unimportant, the accretion process is mediated by an accretion disk. If the magnetic field of the WD is sufficient to prevent the formation of the disk, this gives rise to the so-called polars. If, on the other hand, a disc can form but is truncated by the magnetosphere of the central WD, the systems are classified as intermediate-polars (IPs; see e.g. Warner, 1995, for a review). Disk-mediated systems with mass transfer rates sufficiently low to prevent the disk from remaining persistently in a fully ionized state are classified as Dwarf novae (DNe). DNe are characterized by exhibiting eruptions where the system brightens typically 2-5 magnitudes for a relatively short period of time. Such eruptions are thought to be the consequence of a sudden increase in the mass accretion rate onto the compact object due to the accretion disk transitioning from a neutral to a fully ionized state (e.g., Lasota, 2016).
@@
WZ Sge-systems are highly-evolved short orbital period DNe which exhibit even larger amplitude outbursts (up to $\sim 8-9$ mag). These so-called super-outbursts are characterized by prolonged outburst durations (approximately 30 days), and the appearance of superhumps — low-amplitude variations near the system’s orbital period — shortly after reaching peak brightness (Smak, 1993, 2009; Patterson et al., 2005; Kato, 2015; Hameury, 2020). These super-outbursts are typically followed by a series of short-term rebrightenings or echo outbursts which seem to be characteristic of accreting binaries with extreme mass-ratios. These rebrightenings cannot be explained by the standard disk instability model (DIM; e.g. Hameury, 2020). To reconcile observations with theoretical predictions, Campana et al. (2018) recently proposed an alternative explanation where the luminosity drops during outburst decay signal a temporary transition to a magnetic propeller state. This is an appealing framework as it could help to explain the behaviour of accreting systems at different scales, i.e. X-ray binaries, young stellar objects and CVs. This framework also aligns well with the proposed spin period of 27.87 seconds of WZ Sge (Patterson, 1980; Patterson et al., 1998). However, using HST time-resolved spectroscopy, Georganti et al. (2022) showed that WZ Sge does not exhibit any propeller footprint during the proposed propeller state (c.f. Eracleous & Horne, 1996). Nonetheless, truncation of the inner disk is required to explain the relatively high X-ray luminosity. Against this background, Hameury (2020) proposed that the X-ray excess observed in WZ Sge could (also) be a consequence of evaporation of the inner disk, while the quiescent phases between outbursts are produced by overflows of the mass-transfer stream impacting the inner regions of the disc, similar to what is proposed to explain IW And-systems and some nova-like CVs (Kimura et al., 2020; Castro Segura et al., 2021). Regardless of whether magnetic fields are responsible for the X-ray luminosities in WZ Sge stars, they are often invoked to explain the very long recurrence timescales in these systems (Kato, 2015).
@@ 2 Observations and data reduction
The OPtical TIming CAMera (OPTICAM; Castro et al., 2019, 2024) is a new high-cadence, multi-band camera mounted on the 2.1-meter telescope at the San Pedro Mártir Observatory (OAN-SPM), in México. OPTICAM is equipped with three Andor Zyla 4.2-Plus sCMOS cameras with three $2048\times 2048$ pixels and a set of SDSS filters ($u^{\prime}g^{\prime}r^{\prime}i^{\prime}z^{\prime}$), allowing coverage in the $3200{\rm AA}-1.1{\rm\mu m}$ range. The field of view (FoV), is $\approx 5\times 5~{}{\rm arcmin^{2}}$ with a pixel-scale of $\simeq 0.15^{\prime\prime}/{\rm pix}$.
GOTO 0650 was observed in the $g^{\prime},r^{\prime},i^{\prime}$ bands during the nights of October 26, 28, 30 and November 2nd and 4th of 2024, with a cadence of $3$-s, $3$-s, $5$-s, $15$-s, $15$-s respectively. The typical length of the observations was $\simeq 4.2$ hrs; the details of the observations can be found in Table 1. The data were primarily reduced using version 1.13.0 of the photutils Python package (Bradley et al., 2024). Cosmic rays were clipped from all images using the L.A.Cosmic algorithm (van Dokkum, 2001) as implemented in Astro-SCRAPPY version 1.2.0 (McCully et al., 2018). For each image, two dimensional background images were calculated using the SExtractorBackground estimator from photutils with a box_size of 32 pixels. Image segmentation was used to identify sources in the background and subtracted images using the SourceFinder routine from photutils, which combines source detection and deblending. To identify sources, we set the npixels parameter to 8 pixels, and the threshold parameter to $5\times{\rm background~{}RMS}$. The background RMS was estimated using StdBackgroundRMS. All other parameters were left to their default values.
We used the optimal photometry algorithm described in Naylor (1998) to perform photometry on the background-subtracted images. We model the PSF of OPTICAM’s three cameras using two-dimensional Gaussians for each camera. We stacked the images from each epoch for each camera to get the PSF parameters. We then identified sources in the stacked images using SourceFinder, which assigns a semi-major and semi-minor standard deviation to each source. The PSF for each camera in each epoch was then modelled using the median semi-major and semi-minor standard deviations from the corresponding stacked image. We then used differential photometry to correct for atmospheric variability; relative light curves were computed using the same (non-variable) reference stars in each night.
@@ 3 Analysis and results
The high cadence photometric light curves are presented in the left panels of Figures 2 and 3, the comparison stars in these figures are always the same, so the change in relative flux between epochs is meaningful. As can be seen in Figure 1, epochs 1 and 2 were taken towards the end of the superoutburst, epoch 3 was taken during the decline, epochs 4 and 5 occurred during the onset of the dip between the main outburst structure and the rebrightenings, when the source was $\sim 4-5$ mags fainter than the first observation but $\sim 3$ mags brighter than the quiescent level. During epochs 1–3, the source was dominated by the regular superhumps reported by Tampo (2024), with the dispersion in the light curves reducing as the accretion disk got fainter. In epochs 4 and 5 there is still a slow modulation present with similar frequency as before, as can be seen in Figure 3, the general trend is similar but the $g$-band exhibits higher amplitude (than at least the $r$-band). In both of these epochs, we can see short-term variability in the $g$-band, suggesting the presence of an additional component contributing to the modulation of the light curve.
To characterise the short term variability of GOTO 0650 we have computed the generalized Lomb-Scargle periodogram (LSP; Lomb, 1976; Scargle, 1982; VanderPlas, 2018) for each light curve as implemented in astropy version 6.1.2. In cases where our light curves contained no gaps, we also computed the power spectrum using the fast Fourier transform as implemented in NumPy version 2.0.1. We normalised our power spectra according to:
$P=\frac{2\delta t}{N}|DFT|^{2},$ (1)
where $\delta t$ is the time sampling of the light curve and $N$ is the number of points in the light curve. This normalisation is defined such that the integrated power spectrum yields the variance of the light curve (Vaughan et al., 2003). We normalised our LSPs according to:
@@
where $\delta t$ is the time sampling of the light curve and $P_{\rm LSP}$ represents the unnormalised LSP powers (equivalent to $|DFT|^{2}/N$ for uniformly sampled time series).
To estimate the confidence levels of the proposed signals in our power spectra, we model the power spectrum continuum analytically following the method described in Appendix A of Vaughan (2005). We compute confidence thresholds by scaling our power spectrum model by $-2{\rm ln}(\epsilon/n^{\prime})$, where $\epsilon$ is the desired false alarm probability and $n^{\prime}$ is the number of power spectrum points used to fit the model. The power spectra for each of the light curves is shown in the right panel of Figures 2 and 3. We found a periodicity of $\rm P_{\omega}=148.2$ s with more than $99.99\%$ confidence only in the $g$-band of epoch 4, though we also note a non-statistically significant signal at the same frequency in the r band of the same epoch. Unfortunately, the weather conditions during epoch 5 were far from ideal, with variable cloud coverage; this is reflected in our ability to obtain differential photometry, as seen in the corresponding epoch of Fig 3. In turn, we are not able to recover any signal from the power spectrum but, despite this, the $g$-band lightcurve from that same epoch exhibits short-term flares qualitatively similar to those observed in epoch 4. We derive the analytical false alarm probability of our epoch 4 $g$-band detection to be $1.16\times 10^{-6}$; the probability of detecting a signal of this significance in at least one of our 14 power spectra by chance (i.e., when there is no underlying signal) is 0.0016 per cent.
@@
Figure 3: Same as Fig. 2 but for epochs 4 and 5. For the $g$-band periodogram in epoch 4, the 99.99 per cent confidence threshold is shown with a dotted line.
@@
To determine the centre and $1\sigma$ uncertainty of the frequency detected in epoch 4 we have performed a bootstrap analysis (e.g. Ivezić et al., 2014). In Figure 4, we present the bootstrapped frequency distributions as well as the LSPs for each band in epoch 4. For the LSPs, 99.99 per cent confidence thresholds are shown. These confidence thresholds were estimated by simulating 10,000 white-noise light curves for each camera. Each simulated light curve had the same time sampling as the observed light curve, and fluxes were generated from the following Gaussian $f_{i}\sim\mathcal{N}(\bar{f},\sigma_{f_{i}})$, where $\bar{f}$ represents the mean flux of the observed light curve and $\sigma_{f_{i}}$ represents the uncertainties on the observed flux values. As can be seen in Figure 3, the periodograms of Epoch 4 are white-noise-dominated beyond 5 mHz, and so our estimated confidence thresholds are unlikely to be significantly biased in this frequency range.
@@
Figure 4: Left: frequency distributions inferred from bootstrapped Epoch 4 light curves. Right: Epoch 4 LSPs. In all panels, the dotted vertical line corresponds to a frequency of 6.7 mHz. The horizontal line represents the $99.99\%$ confidence threshold.
As can be seen in Figure 4, the 6.7 mHz signal is persistent in most bootstrapped light curves in the $g$-band. The 6.7 mHz signal is somewhat less persistent in the $r$-band, and even less so in the $i$-band. From these frequency distributions, we find peak frequencies of $6.73\pm 0.02$ mHz ($g$-band), $6.73\pm 0.02$ mHz ($r$-band), and $6.75\pm 0.02$ mHz (i-band). These frequencies correspond to periods of $148.5\pm 0.4$ s, $148.7\pm 0.4$ s, and $148.2\pm 0.5$ s, respectively.
@@
To test the veracity of the signal, we have conducted power spectra of the other stars as well as the background itself, in case the signal was produced by artefacts from the electron noise in the detectors. We have not found any significant signal in any of these. In Figure 5, we present the LSP of the raw counts in the Epoch 4 $g$-band light curve for GOTO 0650 (top panel), along with the LSP of the corresponding local background (bottom panel). There is a strong peak in the periodogram of the raw flux at 6.7 mHz that is not seen in the periodogram of the local background. Since the periodogram in the top panel was computed from the raw flux, this figure shows that the 6.7 mHz signal cannot be attributed to our reference stars. Moreover, the lack of any prominent signal in the local background makes it unlikely that the 6.7 mHz signal is a result of some instrumental/systematic effect.
@@ 4 Discussion
When there is more than one periodicity in a CV, typically, their periodograms exhibit an excess of power due to the interference of these signals, a.k.a. the beat period. We can use the presence of “regular” superhumps to estimate the orbital period. A period of $\rm P_{sh}=91.5$ min places GOTO 0650 below the period gap of CVs as its orbital period would be 10% shorter than $\rm P_{sh}$ (e.g. Kato et al., 2008). Assuming these values we can estimate the beat frequency as $\rm\nu_{beat}=\nu_{orb}-\nu_{sh}$ (where $\rm\nu_{beat},\ \nu_{orb}\ and\ \nu_{sh}$ are the beat, orbital and superhump frequency respectively), this results in a beat frequency $\rm\nu_{beat}\simeq 20~{}\mu Hz$, $\sim 300$ times slower than the signal we observe and more than a factor of $\sim 30$ from other WZ Sge-IPs (Woudt et al., 2012). Dwarf nova oscillations (DNOs) have been proposed to explain some variable periodicity in outbursting CVs (e.g. Marsh & Horne, 1998; Woudt & Warner, 2002), these are vertically extended regions of the disk or blobs being irradiated by the hot WD and boundary layer, which produce a signal corresponding to their Keplerian orbit. These signals are typically of the order of 10 seconds and evolve to slightly lower frequencies as the outburst evolves (up to 40s). We do not detect any significant signal during the main outburst, therefore we discard this scenario. The accretion-induced heating from the outburst makes non-radial pulsations of the WD also unlikely since the temperature of the WD should be too hot to be in the instability strip (c.f. Clemens, 1993; Toloza et al., 2016, but also see Szkody 2021), given its pre-outburst temperature (Killestein et al. submitted). Finally, Veresvarska et al. (2024) proposed a precessing inner disk producing QPOs in a handful of CVs, however, these observed frequencies are typically one order of magnitude slower. Therefore we conclude that GOTO 0650 is likely a WZ Sge-IP, but, the stability of the signal would need to be tested once the system is back to quiescent levels.
@@
Under this assumption, we can estimate the magnetic field assuming canonical values for typical WZ Sge stars from Knigge et al. (2011), i.e. $\dot{\rm M}_{tr}\simeq 10^{-11}~{}\mathrm{M_{\odot}\,yr^{-1}}$, and a typical $0.8M_{\odot}$ WD (Zorotovic et al., 2011; Pala et al., 2017). Assuming the magnetosphere of the white dwarf truncates the disk at the co-rotation radius, i.e. $\rm r_{M}=\xi r_{A}=r_{CO}$ (with $\rm\xi\simeq 0.5$ e.g. Ghosh & Lamb, 1979; Long et al., 2005), we can estimate the the magnetic field, $\rm B$, to be $\sim 1-5\times 10^{4}~{}{\rm G}$ for $\rm P_{spin}=1-2\times P_{\omega}$ and the $\dot{\rm M}_{acc}$ required to push the magnetosphere to the surface of the WD would be $\rm 6\times 10^{-9}~{}M_{\odot}~{}yr^{-1}$ and $\rm 3\times 10^{-8}~{}M_{\odot}~{}yr^{-1}$ for $\rm P_{\omega}=1$ and $2\times P_{spin}$ respectively. Given that we only detect significant periodicity at low luminosities $\rm L\lesssim 10\times L_{quiescence}$ and the accretion luminosity is proportional to the mass accretion rate (${\rm L}_{acc}\propto\dot{\rm M}_{acc}$), we suggest that $\rm P_{\omega}$ is the fundamental frequency of the spin period. If confirmed, GOTO 0650 would not only be another example of how magnetic fields can be dynamically important in highly evolved CVs but also would bridge the gap between intermediate- and fast- rotating AWDs.

### v2

@@ 1 Introduction
Cataclysmic variable stars (CVs) are binary systems in which a low-mass secondary star, filling its Roche lobe, transfers mass to a white dwarf (WD) primary. In systems where the magnetic field is dynamically unimportant, the accretion process is mediated by an accretion disc. If the magnetic field of the WD is sufficient to prevent the formation of the disc, this gives rise to the so-called polars. If, on the other hand, a disc can form but is truncated by the magnetosphere of the central WD, the systems are traditionally classified as intermediate-polars (IPs), however, recent findings suggest that some IPs may be disc-less (e.g. Littlefield et al., 2021). Disc-mediated systems with mass transfer rates sufficiently low to prevent the disc from remaining persistently in a fully ionized state are classified as Dwarf novae (DNe; see e.g. Warner, 1995, for a review on CVs). DNe are characterized by exhibiting eruptions where the system brightens typically 2-5 magnitudes for a relatively short period of time. Such eruptions are thought to be the consequence of a sudden increase in the mass accretion rate onto the compact object due to the accretion disc transitioning from a neutral to a fully ionized state (e.g., Lasota, 2016).
@@
WZ Sge-systems are highly-evolved short orbital period DNe which exhibit even larger amplitude outbursts (up to $\sim 8-9$ mag). These so-called super-outbursts are characterized by prolonged outburst durations (approximately 30 days), and the appearance of superhumps — low-amplitude variations near the system’s orbital period — shortly after reaching peak brightness (Smak, 1993, 2009; Patterson et al., 2005; Kato, 2015; Hameury, 2020). These super-outbursts are typically followed by a series of short-term rebrightenings or echo outbursts which seem to be characteristic of accreting binaries with extreme mass-ratios. These rebrightenings cannot be explained by the standard disc instability model (DIM; e.g. Hameury, 2020). To reconcile observations with theoretical predictions, Campana et al. (2018) recently proposed an alternative explanation where the luminosity drops during outburst decay signal a temporary transition to a magnetic propeller state. This is an appealing framework as it could help to explain the behaviour of accreting systems at different scales, i.e. X-ray binaries, young stellar objects and CVs. This framework also aligns well with the proposed spin period of 27.87 seconds of WZ Sge (Patterson, 1980; Patterson et al., 1998). However, using HST time-resolved spectroscopy, Georganti et al. (2022) showed that WZ Sge does not exhibit any propeller footprint during the proposed propeller state (c.f. Eracleous & Horne, 1996). Nonetheless, truncation of the inner disc is required to explain the relatively high X-ray luminosity. Against this background, Hameury (2020) proposed that the X-ray excess observed in WZ Sge could (also) be a consequence of evaporation of the inner disc, while the quiescent phases between outbursts are produced by overflows of the mass-transfer stream impacting the inner regions of the disc, similar to what is proposed to explain IW And-systems and some nova-like CVs (Kimura et al., 2020; Castro Segura et al., 2021). Regardless of whether magnetic fields are responsible for the X-ray luminosities in WZ Sge stars, they are often invoked to explain the very long recurrence timescales in these systems (Kato, 2015).
@@ 2 Observations and data reduction
The OPtical TIming CAMera (OPTICAM; Castro et al., 2019, 2024) is a new high-cadence, multi-band camera mounted on the 2.1-meter telescope at the San Pedro Mártir Observatory (OAN-SPM), in México. OPTICAM is equipped with three Andor Zyla 4.2-Plus sCMOS cameras with three $2048\times 2048$ pixels and a set of SDSS filters ($ugriz$), allowing coverage in the $3200{\rm AA}-1.1{\rm\mu m}$ range. The field of view (FoV), is $\approx 5\times 5~{}{\rm arcmin^{2}}$ with a pixel-scale of $\simeq 0.15^{\prime\prime}/{\rm pix}$.
GOTO 0650 was observed in the $g,r,i$ bands during the nights of October 26, 28, 30 and November 2nd and 4th of 2024. The sky conditions were photometric for all the nights except for the last one, during which variable high clouds were present; this is reflected in our ability to recover the i band photometry in epoch 5. The details of the observations can be found in Table 1. The data were primarily reduced using version 1.13.0 of the photutils Python package (Bradley et al., 2024). Cosmic rays were clipped from all images using the L.A.Cosmic algorithm (van Dokkum, 2001) as implemented in Astro-SCRAPPY version 1.2.0 (McCully et al., 2018). For each image, two dimensional background images were calculated using the SExtractorBackground estimator from photutils with a box_size of 64 pixels for $2\times 2$ binning or 42 pixels for $3\times 3$ binning. Image segmentation was used to identify sources in the background and subtracted images using the SourceFinder routine from photutils, which combines source detection and deblending. To identify sources, we set the npixels parameter to 32 pixels for $2\times 2$ binning and 14 pixels for $3\times 3$ binning. The threshold parameter was set to $5\times{\rm background~{}RMS}$, and the background RMS was estimated using StdBackgroundRMS. All other parameters were left to their default values.
We used the optimal photometry algorithm described in Naylor (1998) to perform photometry on the background-subtracted images. We model the PSF of OPTICAM’s three cameras using two-dimensional Gaussians for each camera. We stacked the images from each epoch for each camera to get the PSF parameters. We then identified sources in the stacked images using SourceFinder, which assigns a semi-major and semi-minor standard deviation to each source. The PSF for each camera in each epoch was then modelled using the median semi-major and semi-minor standard deviations from the corresponding stacked image. We then used differential photometry to correct for atmospheric variability; relative light curves were computed using the summed fluxes of the same (non-variable) reference stars in each night.
@@ 3 Analysis and results
The high cadence photometric light curves are presented in the left panels of Figures 2 and 3, the comparison stars in these figures are always the same (namely Gaia EDR3 1003219393009879424 and 1003225264228638336), so the change in relative flux between epochs is meaningful. As can be seen in Figure 1, epochs 1 and 2 were taken towards the end of the superoutburst, epoch 3 was taken during the decline, epochs 4 and 5 occurred during the dip between the main outburst structure and the rebrightenings, when the source was $\sim 4-5$ mags fainter than the first observation but $\sim 3$ mags brighter than the quiescent level. During epochs 1–3, the source was dominated by the regular superhumps reported by Tampo (2024), with the dispersion in the light curves reducing as the accretion disc got fainter. In epochs 4 and 5 there is still a slow modulation present with similar frequency as before, as can be seen in Figure 3, the general trend is similar but the $g$-band exhibits higher amplitude (than at least the $r$-band). In both of these epochs, we can see short-term variability in the $g$-band, suggesting the presence of an additional component contributing to the modulation of the light curve.
To characterise the short term variability of GOTO 0650 we have computed the generalized Lomb-Scargle periodogram (LSP; Lomb, 1976; Scargle, 1982; VanderPlas, 2018a) for each light curve as implemented in astropy version 6.1.2. We normalised our LSPs according to:
@@
where $\delta t$ is the time resolution of the light curve and $P_{\rm LSP}$ represents the unnormalised LSP powers. This normalisation is defined such that the integrated periodogram yields the variance of the light curve [^1] Provided the time series is uniformly sampled and the LSP is evaluated at the Fourier frequencies. (Vaughan et al., 2003).
To estimate the confidence levels of the proposed signals in our LSPs, we model the continuum analytically following the method described in Appendix A of Vaughan (2005), which we note is only applicable to the LSP if the time series is uniformly sampled and the LSP is evaluated at the Fourier frequencies (VanderPlas, 2018b). We compute confidence thresholds by scaling our continuum model by $-2{\rm ln}(\epsilon/n^{\prime})$, where $\epsilon$ is the desired false alarm probability and $n^{\prime}$ is the number of LSP ordinates used to fit the model. The LSP for each of the light curves is shown in the right panel of Figures 2 and 3. We found a periodicity of $\rm P_{\omega}=148.2$ s with more than $99.99$ per cent confidence only in the $g$-band of epoch 4, though we also note a non-statistically significant signal at the same frequency in the r band of the same epoch (72.25 per cent confidence). In turn, we are not able to recover any signal from the LSP but, despite this, the $g$-band lightcurve from that same epoch exhibits short-term flares qualitatively similar to those observed in epoch 4. We derive the analytical false alarm probability of our epoch 4 $g$-band detection to be $1.16\times 10^{-6}$; the probability of detecting a signal of this significance in at least one of our 14 LSPs by chance (i.e., when there is no underlying signal) is 0.0016 per cent.
@@
Figure 3: Same as Fig. 2 but for epochs 4 and 5. For the $g$-band periodogram in epoch 4, the 99.99 per cent confidence threshold is shown with a dotted line. The target was not detected in the $i$ band during epoch 5, therefore, only g and r bands are shown.
To determine the centre and $1\sigma$ uncertainty of the frequency detected in epoch 4, and the coherence of this modulation, we performed a bootstrap analysis (e.g. Ivezić et al., 2014). We note that bootstrapping can also be used to reject sampling aliases (e.g., Southworth et al., 2006), which may be present in the $i$-band LSP. From this analysis, we found peak frequencies of $6.73\pm 0.02$ mHz ($g$-band), $6.73\pm 0.02$ mHz ($r$-band), and $6.75\pm 0.02$ mHz (i-band). These frequencies correspond to periods of $148.5\pm 0.4$ s, $148.7\pm 0.4$ s, and $148.2\pm 0.5$ s, respectively. To further test the coherence, we fit a Lorentzian to the Epoch 4 $g$-band periodogram (Belloni et al., 2002) on top of our noise model (Vaughan, 2005, top right panel in Figure 3) and found a Q factor of 3422. The bootstrapped uncertainties and high Q factor are both indicative of a highly coherent signal.
@@
To visualise the modulations identified in Epoch 4, we phase folded these light curves on the peak frequencies inferred via our bootstrap analysis; we present these phase folded light curves in Figure 4, along with localised LSPs. Prominent modulations are only apparent in the g-band, with a peak-to-peak amplitude of roughly 10 per cent (compared to an average $1\sigma$ white noise amplitude of 2.6 per cent). The $3\sigma$ upper limits using the average white noise amplitude for the Epochs 1, 2, 3 and 5 in $g$-band is 4.8, 4.5, 14.5 and 10.5 per cent respectively. For the LSPs, 99.99 per cent confidence thresholds were estimated by simulating 10,000 white-noise light curves for each band. Each simulated light curve had the same time sampling as the observed light curve, and fluxes were generated from the following Gaussian $f_{i}\sim\mathcal{N}(\bar{f},\sigma_{f_{i}})$, where $\bar{f}$ represents the mean flux of the observed light curve and $\sigma_{f_{i}}$ represents the uncertainties on the observed flux values. As can be seen in Figure 3, the periodograms of Epoch 4 are white-noise-dominated beyond 5 mHz, and so our estimated confidence thresholds are unlikely to be significantly biased in this frequency range.
@@
Figure 4: Left: Epoch 4 light curves folded on a period of 148.6 s (corresponding to 6.7 mHz), the black line represents the median values for the phase binned folded light curve, the zero phase is arbitrary but common across bands. Right: Epoch 4 LSPs. In the LSPs, the dotted vertical line corresponds to a frequency of 6.7 mHz and the horizontal line represents the $99.99$ per cent confidence threshold.
@@
To test the veracity of the signal, we have computed LSPs of the other stars as well as the background itself, in case the signal was produced by artefacts from the electron noise in the detectors. We have not found any significant signal in any of these. In Figure 5, we present the LSP of the raw Epoch 4 $g$-band light curve for GOTO 0650 (top panel), along with the LSP of the corresponding local background (bottom panel). Figure 5 shows a prominent peak in the periodogram of the raw flux at 6.7 mHz that is not seen in the periodogram of the local background. This rules out the 6.7 mHz signal being attributable to our reference stars, or a result of some instrumental/systematic effect, and instead suggests that the signal is intrinsic to GOTO 0650.
@@ 4 Discussion
Dwarf nova oscillations (DNOs) have been proposed to explain some variable periodicity in outbursting CVs (e.g. Marsh & Horne, 1998; Woudt & Warner, 2002), these are vertically extended regions of the disc or blobs being irradiated by the hot WD and boundary layer, which produce a signal corresponding to their Keplerian orbit. These signals are typically of the order of 10 seconds and evolve to slightly lower frequencies as the outburst evolves (up to 40s). We do not detect any significant signal during the main outburst, therefore we discard this scenario. The accretion-induced heating from the outburst makes non-radial pulsations of the WD also unlikely since the temperature of the WD should be too hot to be in the instability strip (c.f. Clemens, 1993; Toloza et al., 2016, but also see Szkody 2021), given its pre-outburst temperature (Killestein et al., 2025). Finally, Veresvarska et al. (2024) proposed a precessing inner disc producing QPOs in a handful of CVs, however, these observed frequencies are typically one order of magnitude slower. Therefore we conclude that GOTO 0650 is likely a WZ Sge-IP, but, the stability of the signal would need to be tested once the system is back to quiescent levels.
@@
The data from Epoch 4 was obtained 2-3 days after the decline from the main outburst; therefore, the disc and magnetic torques would have been far from equilibrium during this epoch, making any estimation of the magnetic field highly uncertain. However, we can set a lower limit on the magnetic field as the source’s magnetospheric radius must be larger than the WD radius. To do this, we need to estimate the mass accretion rate; for a canonical mass transfer rate of $\dot{\rm M}_{tr}\simeq 10^{-11}~{}\mathrm{M_{\odot}\,yr^{-1}}$ in WZ Sge stars (Knigge et al., 2011), typically expending $\gtrsim 30$ years in quiescence between outbursts, and with an outburst duration of $\simeq 30$ d, we estimate the average mass accretion rate during the outburst to be $<\dot{M}_{acc}>\simeq 7\times 10^{-9}~{}\mathrm{M_{\odot}\,yr^{-1}}$ in a conservative mass transfer scenario. Alternatively we can obtain the peak mass accretion rate via the peak absolute magnitude $\rm M_{V}$ vs orbital period ($P_{\rm orb}$) and its relation with $\dot{\rm M}_{acc}$ (e.g. Warner, 1987; Patterson, 2011). For a source close to the period minimum (as GOTO 0650 is), we find the absolute magnitude at (super outburst) peak to be $\rm M_{V}\simeq 5$, consistent with the value derived by Killestein et al. (2025). In turn, this value give us an estimation of the mass accretion rate of $\dot{\rm M}_{acc}\simeq 2\times 10^{-8}~{}\mathrm{M_{\odot}\,yr^{-1}}$ (e.g. Warner, 1987). We can therefore assume $\dot{\rm M}_{acc}\sim 10^{-8}~{}\mathrm{M_{\odot}\,yr^{-1}}$ a reasonable mass accretion rate during the main outburst of GOTO 0650. During Epoch 4 the source was 2 orders of magnitude fainter, therefore, assuming the bulk observed luminosity during this epoch being dominated by the accretion disc and given that this scales with the mass accretion rate (${\rm L}_{acc}\propto\dot{\rm M}_{acc}$), we can adopt $\dot{\rm M}_{acc}\sim 10^{-10}~{}\mathrm{M_{\odot}\,yr^{-1}}$ during epoch 4 as a reasonable approximate reference value for our calculations (acknowledging that it excludes corrections or additional spectral components; e.g., the hot spot). If the magnetosphere of the white dwarf truncates the disc at the co-rotation radius, i.e. $\rm r_{M}=\xi r_{A}=r_{CO}$ (with $\rm\xi\simeq 0.5$ e.g. Ghosh & Lamb, 1979; Long et al., 2005), we can set a lower limit of the magnetic field, $\rm B$, at which $\rm r_{M}=R_{WD}$. For a typical $0.8~{}{\rm M_{\odot}}$ WD in a CV (Zorotovic et al., 2011; Pala et al., 2017). this would give us a lower limit of $\rm B\sim 2\times 10^{4}~{}G$. This argument cannot be used for Epoch 3, since the disc is experiencing a rapid change which will keep it even further from a stationary configuration. The same is not true for Epoch 2, thus allowing us to set an upper limit in the magnetic field strength. Assuming the disc “pushed” magnetosphere to the surface of the WD during the outburst, we can constrain the magnetic field to be $B\lesssim 4\times 10^{4}~{}G$.
In the hypothetical case of the disc and magnetic torques being in equilibrium during Epoch 4, we could estimate the magnetic field of the system, for $\rm P_{\omega}=1$ and $2\times P_{spin}$ this would correspond to a magnetic field of $\rm 5\times 10^{4}~{}G$ and $\rm 10^{5}~{}G$ respectively. The latter of these values would require a $\dot{\rm M}_{acc}\simeq 10^{-7}~{}\mathrm{M_{\odot}\,yr^{-1}}$ to push the magnetosphere down to the WD surface, this scenario not only goes against our constraints from above but it would make the system enter a super-Eddington wind regime (c.f. Ma et al., 2013). We therefore suggest that $\rm P_{\omega}$ may be the fundamental frequency of the spin period, and the WD’s magnetic would be of the order of $B\sim 10^{4}~{}G$. However, quiescent observations are required to test this hypothesis. If confirmed, GOTO 0650 would not only be another example of how magnetic fields can be dynamically important in highly evolved CVs but also would bridge the gap between intermediate- and fast-rotating AWDs.